The current induced in 100 ohm coil when the magnetic flux decreases from 1 Wb to 0.1 Wb in 0.1 s,  is

Grade:12

1 Answers

Lokesh Rekha
5 years ago
as rate of change of flux is emf. 
Therefore, average emi induced is equal to 1-0.1/0.1 = 9 volts.
Current induced = 9/100 = 0.09 amperes